OpenStudy (anonymous):

Convert each measure to degrees. Round answers to the nearest hundredth if needed.

OpenStudy (anonymous):

\[a.) \frac{ \pi }{ 9 } b.) \frac{ -2\pi }{ 3 }\]

OpenStudy (anonymous):

@hartnn :D

hartnn (hartnn):

to convert radians into degrees, multiply it by 180/pi example : pi/4 radians = pi/4 * 180/pi = 180/4 = 45 degrees. you try both

hartnn (hartnn):

note : everytime pi will get cancelled.

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

pi/9 radians = pi/9 * 180/pi = 180/9 = 20 degrees.

OpenStudy (anonymous):

is that right?

hartnn (hartnn):

absolutely correct! :D

OpenStudy (anonymous):

omg :D k next one :D

OpenStudy (anonymous):

-2pi/3 radians = -2pi/3 * 180/-2pi = -90/3 = -30 degrees.

hartnn (hartnn):

how u got 90/3 ? there is no -2 in the denominator, its 3

hartnn (hartnn):

-2 * (180/3) =...

OpenStudy (anonymous):

ohh its *

hartnn (hartnn):

the other (easier) way to look at this will be 'just replace pi with 180' because pi radians = 1801 degrees.

OpenStudy (anonymous):

lol so its -120 degrees

hartnn (hartnn):

** pi radians = 180 degrees.

OpenStudy (anonymous):

can you give me the equation for b please sorry I'm just confused in the steps now

hartnn (hartnn):

-120 degrees is correct.

OpenStudy (anonymous):

ohh ok :) thank you so much :)

hartnn (hartnn):

welcome ^_^